a tennis ball is served horizontally from 2.4 above the ground at 30 m/s the net is 12 m away and 0.9m high will be clear the net? a. 0.64m b. 1.62m c 2.72m d.none ans is optn B

Grade:12th pass

1 Answers

Arun
5 years ago
The time for the ball to reach the net:
distance = speed x time
t = 12/30
t = 0.4 seconds

a) Height of ball after 0.4 seconds:
h = 2.4 - ut - 1/2 at²; u = 0
h = 2.4 - 1/2 x 9.81 x (0.4)²
h = 1.62 m